# Training Domain Draft Models for Speculative Decoding: Best Practices and Insights

### v1

@@ 4 Results and Discussion
Offline vs. Online Distillation Our experiments show that offline distillation consistently outperforms online distillation across all three domains (see Figure 2), and this trend holds across different dataset sizes (see Figure 3). A key insight into this trend is that offline distillation leverages supervision from all completion tokens, providing richer learning signals for the draft model. Furthermore, as training data increases, the offline approach gains a growing advantage over online approach in Biology and Chinese. In Biology, offline surpasses online by 11.4% to 24.7% in acceptance rate as data expands from 2k to 19k; in Chinese, the gap widens from 11.2% to 14.1%. We also find out that for training with in-domain user queries data (Data Scenario I), offline distillation can benefit from higher learning rate while online distillation requires a lower learning rate, as detailed in Appendix D.1. However, for Data Scenario II and III where the training data exhibits domain shifts from evaluation data, offline distillation would also prefer a lower learning rate (see Table  2 and 3).
@@
Effectiveness of Synthetic Data (Magpie) We examine the effectiveness of Magpie synthetic data for aligning draft models in Biology and Chinese domains. While not as effective as training on in-domain data, synthetic data still yields meaningful improvements (see Figure 2), making it a viable approach for training an initial draft model before target model deployment. Notably, the performance gap between user query finetuning and synthetic data finetuning is larger in Biology than in Chinese. This can be attributed to the fact that the biomedical target model primarily generating diagnostic and medical-related completions, which exhibit a greater domain shift from real user queries that span biological topics. We also find that offline training with Magpie data benefits from a smaller learning rate than in-domain user query data and achieves over 90% of the acceptance rate observed with in-domain training (see 2). This is likely because Magpie data exhibits some domain shift from the evaluation data, and thus a larger LR leads to overfitting.
Data Size Avg. accept. rate
2k 39.8
4k 41.4
Table 4: Accept. rate(%) of 1B SFT-trained draft model for decoding 70B target model in Biology domain. The baseline is 33.0%.
Scaling to Larger Target Model We extend our experiments to 1B draft and 70B target model (Llama3-OpenBioLLM-70B  [^4] https://huggingface.co/aaditya/Llama3-OpenBioLLM-70B) in Biology domain. As shown in Table 4, SFT is able to improve the draft model’s acceptance rate by 20.5% and 25.4% with 2k and 4k training data respectively.
Training on Related Domain Data When user queries are unavailable, training on related domain data also effectively improves draft model alignment. In Function Calling domain, training on APIGen data [^5] https://huggingface.co/datasets/argilla/apigen-function-calling enhances the draft model’s performance on Hermes-FC [^6] https://huggingface.co/datasets/NousResearch/hermes-function-calling-v1 (Table 3), which serves as a proxy for real user queries. However, domain shift between collected data and user queries necessitates significantly more training samples to achieve comparable results and more careful learning rate selection.
@@ C.1 Table  1
For each line in Table 1, we use a small generic model as the draft model and compare its acceptance rates when decoding a general target model and a domain target model of the same size. The choice of models and evaluation domain datasets are specified in 6. We randomly select 1000 prompts from each dataset as the evaluation set.

### v2

@@ 4 Results and Discussion
Offline vs. Online Distillation Our experiments show that offline distillation consistently outperforms online distillation across all three domains (see Figure 2), and this trend holds across different dataset sizes (see Figure 3). A key insight into this trend is that offline distillation leverages supervision from all completion tokens, providing richer learning signals for the draft model. Furthermore, as training data increases, the offline approach gains a growing advantage over online approach in Biology and Chinese. In Biology, offline surpasses online by 11.4% to 24.7% in acceptance rate as data expands from 2k to 19k; in Chinese, the gap widens from 11.2% to 14.1%. We also find out that for training with in-domain user queries data (Data Scenario I), offline distillation can benefit from higher learning rate while online distillation requires a lower learning rate, as detailed in Appendix D.1. However, for Data Scenario II and III where the training data exhibits domain shifts from evaluation data, offline distillation would also prefer a lower learning rate (see Table  4 and 5).
Target Model Draft Model Biology
Llama3-OpenBioLLM-70B [^2] https://huggingface.co/aaditya/Llama3-OpenBioLLM-70B Llama3.2-1B [^3] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct 33.0%
+ SFT (Data Scenario I) +8.4%
Table 2: Accept. rate(%) of 1B SFT-trained draft model for decoding 70B target model in Biology domain. Training data size is 4k.
Target Model Draft Model Math Coding General Avg.
DS-R1-Qwen-32B [^4] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-32B&qwen2.5-0.5B-Instruct Qwen2.5-0.5B [^5] https://huggingface.co/Qwen/Qwen2.5-0.5B-Instruct 34.7% 35.6% 34.6% 35.0%
+ SFT (Data Scenario II) +5.0% +5.0% +5.0% +5.0%
DS-R1-Llama-70B [^6] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-70B Llama3.2-1B [^7] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct 41.6% 32.3% 26.9% 33.6%
+ SFT (Data Scenario II) +10.0% +17.7% +6.0% +11.2%
Table 3: Accept. rate(%) of SFT-trained draft models for decoding 70B and 32B target models in Reasoning domain. Training data size is 200k.
@@
Effectiveness of Synthetic Data (Magpie) We examine the effectiveness of Magpie synthetic data for aligning draft models in Biology and Chinese domains. While not as effective as training on in-domain data, synthetic data still yields meaningful improvements (see Figure 2), making it a viable approach for training an initial draft model before target model deployment. Notably, the performance gap between user query finetuning and synthetic data finetuning is larger in Biology than in Chinese. This can be attributed to the fact that the biomedical target model primarily generating diagnostic and medical-related completions, which exhibit a greater domain shift from real user queries that span biological topics. We also find that offline training with Magpie data benefits from a smaller learning rate than in-domain user query data and achieves over 90% of the acceptance rate observed with in-domain training (see Table 4). This is likely because Magpie data exhibits some domain shift from the evaluation data, and thus a larger LR leads to overfitting.
Scaling to Larger Target Model In order to assess the effectiveness of training draft models for larger target models, we extend our experiments to models exceeding 8B parameters in the domains of Biology and Reasoning. As presented in Table 2, within the Biology domain, SFT with 4,000 training samples improves the acceptance rate of the 1B draft model by 25.4% for the 70B target model (Llama3-OpenBioLLM-70B).
Additionally, we evaluate our method for two reasoning models distilled from DeepSeek-R1 (DeepSeek-AI et al., 2025), which incorporate extended reasoning processes in their outputs and demonstrate strong performance in reasoning-intensive tasks. To conduct our experiments within the framework of Data Scenario II, we use a multi-domain instruction dataset for training and reasoning-intensive domain datasets for evaluation. Specifically, we sample 200,000 prompts from a combination of the training dataset introduced in Jain et al. (2024) and the Magpie-500K dataset [^8] https://huggingface.co/datasets/Magpie-Align/Magpie-Llama-3.1-Pro-500K-Filtered, covering math, coding, and additional domains. For evaluation, we construct domain-specific datasets for math, coding, and general reasoning by sampling 500 prompts from AIME [^9] https://huggingface.co/datasets/di-zhang-fdu/AIME_1983_2024, BigCodeBench [^10] https://huggingface.co/datasets/bigcode/bigcodebench, and AlpacaEval (Dubois et al., 2024), respectively. As summarized in Table 3, the SFT-trained Qwen 2.5 0.5B Instruct model achieves a 14.3% improvement in acceptance rate when serving as a draft model for the 32B reasoning model. Similarly, the SFT-trained LLaMA 3.2 1B Instruct model exhibits a 33.3% improvement when used as a draft model for the 70B reasoning model. These results indicate that the effectiveness of SFT extends to larger target models, and training on collected domain-relevant data can significantly enhance the alignment of draft models.
@@
Training on Related Domain Data For Data Scenario II, when user queries are unavailable, training on related domain data also effectively improves draft model alignment. In Function Calling domain, training on APIGen data [^13] https://huggingface.co/datasets/argilla/apigen-function-calling enhances the draft model’s performance on Hermes-FC [^14] https://huggingface.co/datasets/NousResearch/hermes-function-calling-v1 (Table 5), which serves as a proxy for real user queries. However, domain shift between collected data and user queries necessitates significantly more training samples to achieve comparable results and more careful learning rate selection. Beyond the Function Calling domain, as discussed in the previous section, SFT on collected domain datasets also greatly improves the performance of the draft models in math, coding, and general reasoning tasks (see Table 3).
@@ C.1 Table  1
For each line in Table 1, we use a small generic model as the draft model and compare its acceptance rates when decoding a general target model and a domain target model of the same size. The choice of models and evaluation domain datasets are specified in 7. We randomly select 1000 prompts from each dataset as the evaluation set.